Contribution Margin and Contribution Margin Ratio

For a recent year, Wicker Company-owned restaurants had the following sales and expenses (in millions):

Sales $16,300
Food and packaging $5,046
Payroll 4,100
Occupancy (rent, depreciation, etc.) 4,264
General, selling, and administrative expenses 2,400
$15,810
Income from operations $490

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What is Wicker Company's contribution margin? Round to the nearest million. (Give answer in millions of dollars.)
_________ $ million

b. What is Wicker Company's contribution margin ratio? Round to one decimal place.
___________ %

c. How much would income from operations increase if same-store sales increased by $1,000 million for the coming year, with no change in the contribution margin ratio or fixed costs? Round your answer to the closest million.
____________ $ million
Answer #1

Variable cost = 5046+4100+(2400*40%) = 10106

a) Contribution margin = 16300-10106 = 6194 million

b) Contribution margin ratio = 6194*100/16300 = 38%

c) Income from operation increase by = 1000*38% = 380 million